06/30/04 5:00 12m The 12 meter telescope was shutdown this morning for summer shutdown. News
06/28/04 8:00 SMT A lightning caused fire has broken out on Mt. Graham about 5.6 miles away from the MGIO site. It has been designated the NUTTALL fire due to it's location close to Nuttall Canyon. Here is a link to the Forest Service web site about this and other fires in the Southwest. News
06/25/04 15:00 ARO The new ARO web server is now online. If you experience any problems, please feel free to email me at tfolkers{at}as{dot}arizona{dot}edu News
06/21/04 8:00 ARO The new replacement web server for the ARO offices has arrived and is being configured. It will be installed shortly and when it does we may experience web outages for brief periods of time. More news on this front later. News
06/20/04 16:45 12m The 23mmsis receiver is cold and operational. Fixed
06/20/04 12:00 SMT The Steward Observatory's Advanced Teen Astronomy Camp took a tour of the Mt. Graham facilities today. Here are some pictures from that field trip. News
06/18/04 16:00 12m The 23mmsis receiver has been warmed up to replace a bad amplifier in the 3mmlo chan #1. Ongoing
06/18/04 18:00 12m The 1mmsis receiver has been shutdown for the season. It just could not stay cold. This problem with warm-ups in the latter parts of the observing season will be addressed this summer. News
06/16/04 8:00 SMT Summer shutdown in underway on Mt. Graham and already several things have been fixed. A new hot water heater for the Hilton was installed. A complete list for the SMT can be found here. News
06/15/04 8:00 SMT The National Forest Service has increased its fire restrictions effective today. Here is the news release in PDF format. News
06/14/04 14:00 12m Already, the new computers for the 12 Meter have begun arriving and are being configured and readied for their installation after the June 30th shutdown. Here is a brief description of the pending process of control system conversion. Here is a complete list of activities for this summers 12 Meter shutdown. News
06/14/04 8:00 ARO The summer shutdown projects and maintenance items are now online here. News
06/10/04 14:00 12m The 12 Meter telescope will be shutdown for summer shutdown at 5:00 MST on the morning of June 30th. News
06/10/04 14:00 SMT The SMT 10 Meter telescope will be shutdown for summer shutdown at 8:00 MST on the morning of June 11th. News
06/09/04 14:00 SMT The new Forbes Filterbank tests were a smashing success and now we are shifting to the construction phase. Here is a (PS) spectral line plot of a 12CO(2-1) scan on IRC+10216, or if you prefer, (PDF). Here is a 12CO(2-1) OTF map of CW Leo taken with the filters. Here is a similarly sampled AOS map for comparison. News
06/07/04 8:00 SMT The new 2048 Channel filter bank system for the SMT is being taken to the telescope for the last prototype testing. In it's prototype configuration, it only contains (48) 1 MHz channels. See here for a brief description. News
06/03/04 12:00 SMT Tests continue using the new 2mmjt. Ongoing
06/02/04 12:00 SMT The 2mm JT receiver is mounted on the right flange and we are beginning system test. News
06/02/04 20:00 12m The frequent reboots of modelo have been traced to a faulty power strip. It has been replaced and all seems well for now. Fixed
06/01/04 14:00 SMT We were able to resurrect the 4-color bolometer and obtained satisfactory results all the way up to 350Ám. Not bad for the first day of June... News

 Copyright Arizona Radio Observatory.
For problems or questions regarding this web contact [tfolkers{at}email{dot}arizona{dot}edu].
Last updated: 02/09/10.